Find the tension in the horizontal part of the supporting cable, and the compression in the boom, elevated to 25.0°, of the crane that lifts a 8900.N load.

1
Expert's answer
2022-03-03T14:24:21-0500

Answer

the tension in the horizontal part of the supporting cable,

"T=wcot\theta\\=8900*cot25°\\=1.91*10^2N"